# Preserving Task-Relevant Information Under Linear Concept Removal

### v1

@@ Abstract
Modern neural networks often encode unwanted concepts alongside task-relevant information, leading to fairness and interpretability concerns. Existing post-hoc approaches can remove undesired concepts but often degrade useful signals. We introduce SPLICE—Simultaneous Projection for LInear concept removal and Covariance prEservation—which eliminates sensitive concepts from representations while exactly preserving their covariance with a target label. SPLICE achieves this via an oblique projection that “splices out” the unwanted direction yet protects important label correlations. Theoretically, it is the unique solution that removes linear concept predictability and maintains target covariance with minimal embedding distortion. Empirically, SPLICE outperforms baselines on benchmarks such as Bias in Bios and Winobias, removing protected attributes while minimally damaging main-task information.
@@ 1 Introduction
In this paper, we seek to address a key drawback of post-hoc concept-removal methods. Our contribution is to introduce SPLICE, a projection that (similar to LEACE or SAL) prevents any linear classifier from predicting a concept, while also preserving the covariance with a task of interest. Mathematically, we construct an oblique projection that places the covariance between the representations and a protected attribute in its kernel, while maintaining the covariance between the representations and the main-task label in its range.
We prove that if a linear classifier is re-fitted after projection without regularization, any two projections that share the same kernel (i.e., that linearly erase the same subspace) will induce identical loss. In that sense, SPLICE and previous methods such as LEACE (Belrose et al., 2023) (as well as more naive versions that do not explicitly aim to maintain the minimality of the projection) are all equivalent. For causal model interventions aiming to interpret its behavior—a situation where the underlying model is necessarily frozen—we argue that SPLICE may perform a more surgical intervention, akin to minimizing the side effects of erasure on related concepts (e.g., removing gender bias while preserving grammatical gender). Empirically, we show that in a realistic classification setting, SPLICE improves fairness in a highly challenging, imbalanced scenario, and removes stereotypes while maintaining correlated factual information.
@@ 2 Related work
Existing linear concept-removal methods: iterative nullspace projection (INLP, Ravfogel et al. (2020)) trains a linear classifier to predict a concept, and projects embeddings to the nullspace of the parameters of the linear classifier. This is repeated until the concept can no longer be predicted by the linear classifier. Relaxed Linear Adversarial Concept Erasure (RLACE, Ravfogel et al. (2022)) trains an orthogonal projection matrix such that a concept cannot be predicted by a linear classifier. These works were followed up by Least-squares Concept Erasure (LEACE, Belrose et al. (2023)), which ensures that no linear classifier can predict the concept (hereafter referred to as linear guardedness) while minimally altering the embeddings. Spectral attribute removal (SAL, Shao et al. (2023)) projects the embeddings orthogonal to the first $k$ eigenvectors of the covariance matrix $\mathrm{Cov}(\bm{x},\bm{z})$. Mean Projection (MP, (Haghighatkhah et al., 2022)) projects embeddings to the nullspace of the mean difference between embeddings with and without concepts. It is equivalent to SAL when the concept is binary. SAL and MP also guarantee linear guardedness.
Linear concept-removal while preserving task-relevant information: previous work suggests that linear concept-removal methods remove task-relevant information in addition to the concept they seek to remove (Belinkov, 2022; Kumar et al., 2022; Guerner et al., 2025). In response, several alternatives have been proposed to address this issue, all removing a different linear subspace (Dev et al., 2021; Holstege et al., 2024; Bareeva et al., 2024; Shi et al., 2024). However, each of these approaches sacrifices linear guardedness in order to retain more task relevant information. An alternative approach is to explicitly optimize for fairness while maintaining task performance (Shen et al., 2021). However, this approach is more resource-intensive, as it cannot be applied to the frozen representations of a pretrained model. Moreover, because it modifies the original representations, it is unsuitable for scenarios where the intervention is intended to simulate causal experiments on the behavior of a pretrained LM, as discussed in section 4.2.
@@ 3 Theory
3.1 SPLICE: Ensuring linear guardedness while preserving task-relevant covariance
The LEACE projection ensures linear guardedness while minimizing the distortion of the features, but is oblivious of the main task of the model. A small expected norm squared may not optimally preserve information that is actually useful for the task at hand. Suppose now there is a random vector $\bm{y}\in\mathcal{Y}$ that represents task-relevant information. Similar to the concept vector $\bm{z},$ it can be binary, one-hot encoded or continuous. We conjecture that task-relevant information in the features $\bm{x}$ is located in the directions that linearly covariate with $\bm{y}.$ In other words, it is located in the column space of the covariance matrix $\bm{\Sigma}_{\bm{x},\bm{y}}=\mathrm{Cov}(\bm{x},\bm{y}).$; indeed, removing this particular subspace completely prevents linear classification (Belrose et al., 2023).
In order to preserve this task-relevant information, we require the affine transformation $r(\bm{x})=\mathbf{P}\bm{x}+\bm{b}$ not only to produce features that are linearly guarded for $\bm{z},$ but also to leave the covariance between $\bm{x}$ and $\bm{y}$ invariant. For this approach we cast the name SPLICE (Simultaneous Projection for LInear concept removal and Covariance prEservation), which can be seen as an extension of LEACE. It eliminates sensitive concepts from representations while exactly preserving their covariance with a target label. The SPLICE optimization problem is formulated in Theorem 1 and its solution is given by equation 4, which is the main theoretical contribution of this paper.
@@ 3.1 SPLICE: Ensuring linear guardedness while preserving task-relevant covariance
$\mathbf{P}^{\star}_{\mathrm{SPLICE}}=\mathbf{W}^{+}\mathbf{V}(\mathbf{U}^{\mathrm{T}}\mathbf{V})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{W},$ (4)
@@
The proof of Theorem 1 is given in Appendix A.1. Compared to the LEACE optimization problem, the only difference is the additional condition of preserving task-relevant information, $\mathbf{P}\bm{\Sigma}_{\bm{x},\bm{y}}=\bm{\Sigma}_{\bm{x},\bm{y}},$ which we refer to as the range constraint. Similar to LEACE, $\mathbf{P}^{\star}_{\mathrm{SPLICE}}$ is an oblique transformation with kernel $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}}).$ The difference lies in the range, which now contains $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ in order to fulfill the range constraint. Intuitively, the freedom that the LEACE optimization problem gives to the choice of the range is partially used to preserve the task-relevant information, i.e., the covariance between $\bm{x}$ and $\bm{y}.$ The remainder of the freedom is used to minimize the distortion caused by the affine transformation, leading to whitening and unwhitening, similar to LEACE.
We note that, perhaps counter-intuitively, SPLICE is not necessarily equivalent to LEACE if $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}})$ and $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ are orthogonal subspaces. Only if those subspaces are orthogonal after whitening, SPLICE and LEACE are equivalent. In that case the orthogonal projection of LEACE then already contains the task-relevant directions $\mathrm{colsp}(\mathbf{W}\bm{\Sigma}_{\bm{x},\bm{y}})$ in its range. We also note that, similar to LEACE (Belrose et al., 2023), in the case of non-centered data, i.e., $\mathbb{E}[\bm{x}]\neq\mathbf{0},$ the optimal affine transformation requires the addition of a constant $\bm{b}^{\star}_{\mathrm{SPLICE}}=\mathbb{E}[\bm{x}]-\mathbf{P}^{\star}_{\mathrm{SPLICE}}\mathbb{E}[\bm{x}].$ Finally, in Figure 1 we give a visual illustration of the steps of the projection matrix suggested by Theorem 1.
@@ 3.2 Last-layer linear concept removal with re-training
The first use case of SPLICE we consider is linear concept removal applied to the embeddings of a DNN, after which a linear classifier is fitted on the transformed embeddings. This can be useful if it is demanded that a predictive model does not make use of sensitive concepts, like gender or race.
If we compare SPLICE with other concept removal methods that guarantee linear guardedness, namely LEACE and SAL (Shao et al., 2023), we observe that they are all projections with the same kernel $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}}).$ They differ in the choice of the range. Interestingly, we find that the predictions of a linear classifier that is trained without regularization on the transformed embeddings are not affected by this choice of the range. In other words, all concept removal methods that ensure linear guardedness will lead to the same predictions after re-training a linear classifier without regularization. We verify this empirically in Appendix B.1.
This result is formalized in Theorem 3.2, in which we consider training a model $f(\bm{x};\bm{\theta})$ that only depends on the embeddings $\bm{x}$ and parameters $\bm{\theta}$ through their inner product. Examples of such models are linear and logistic regression, the latter typically being used as the linear classifier re-trained on the embeddings. Before fitting, an projection is applied to the embeddings. We consider two projections that have the same kernel, but different ranges. Theorem 3.2 shows that, in the case of a strictly convex loss function without regularization, both fitted models lead to the same predictions.
@@ 3.3 When does changing the range matter?
Theorem 3.2 provides a case where SPLICE will lead to the same predictions as other concept removal methods that ensure linear guardedness (e.g., SAL and LEACE). Here, we identify two practical cases where applying projections with the same kernel and different ranges will typically lead to different predictions.
@@
    When re-training the last layer with regularization: if we include a regularization term (such as $||\bm{\theta}||_{2}$ or $||\bm{\theta}||_{1}$) in our loss function $\mathcal{L}$, then it no longer exclusively depends on the parameters via the inner product $\mathrm{\mathbf{X}}\bm{\theta}.$ This will generally lead to $\bm{x}_{\mathcal{A}}^{\mathrm{T}}\bm{\theta}_{\mathcal{A}}^{*}\neq\bm{x}_{\mathcal{B}}^{\mathrm{T}}\bm{\theta}_{\mathcal{B}}^{*}$ for any two projections with the same kernel and different ranges.
@@
    Since the projections considered in this paper are constructed without knowledge of the parameters, these will typically lie outside the orthogonal complements.
    Note that this use case is relevant for language modeling, where re-training the parameters of the last layer is typically not feasible in terms of computational resources and/or data availability.
For these cases we expect SPLICE to outperform other methods that ensure linear guardedness, because it is explicitly designed to preserve task-relevant information. We empirically investigate this hypothesis in the next section.
@@ 4 Experiments
This section is structured as follows. We start by investigating classification tasks when the last layer of the model is re-fitted with regularization. Then, in Section 4.2, we investigate language modeling, when the last layer of the language model (LLM) is not re-trained post-projection. Finally, in order to qualitatively assess the effect of the different projections, we apply SPLICE to black and white image data in Section 4.3.
Across experiments, we compare SPLICE to two other projections: LEACE and SAL (see Section 2). We focus on these projections since, similar to SPLICE, they ensure linear guardedness with regards to a concept, and only differ in choice of range. Furthermore, LEACE and SAL have been shown to outperform other existing concept-removal methods such as INLP and RLACE.
@@ 4.1 Classification where the last layer is re-trained with regularization
Set-up of the experiment: we seek to investigate the impact of each projection as the correlation between the task of interest $(y_{\mathrm{prof}},y_{\mathrm{tox}})$ and the concept to remove $(z_{\mathrm{gender}},z_{\mathrm{lang}})$ becomes stronger. We expect that as the relationship becomes stronger, the difference between SPLICE and other projections becomes greater. The reason is that stronger correlated labels have covariances with the embeddings that are typically more aligned. Removing the concept is then more likely to also remove information about the task of interest.
@@
Results: the results of the experiments for both datasets are given in Figure 2. We focus on overall accuracy as well as worst-group accuracy. Worst-group accuracy is defined as the lowest accuracy for all combinations of the task and concept. A low worst-group accuracy reflects that a model relies on the correlation between the task and concept in the training data (Sagawa et al., 2020). For both datasets, as the relationship between task and concept becomes stronger, SPLICE outperforms the other projections in both accuracy and worst-group accuracy. As the correlation between the task and concept becomes stronger, SAL and LEACE remove a significant part of $\bm{\Sigma}_{\bm{x},\bm{y}},$ contrary to SPLICE. This is illustrated for the Bias in Bios dataset in Figure 7 in Appendix B.2

### v2

@@ Abstract
Modern neural networks often encode unwanted concepts alongside task-relevant information, leading to fairness and interpretability concerns. Existing post-hoc approaches can remove undesired concepts but often degrade useful signals. We introduce SPLINCE—Simultaneous Projection for LINear concept removal and Covariance prEservation—which eliminates sensitive concepts from representations while exactly preserving their covariance with a target label. SPLINCE achieves this via an oblique projection that “splices out” the unwanted direction yet protects important label correlations. Theoretically, it is the unique solution that removes linear concept predictability and maintains target covariance with minimal embedding distortion. Empirically, SPLINCE outperforms baselines on benchmarks such as Bias in Bios and Winobias, removing protected attributes while minimally damaging main-task information.
@@ 1 Introduction
In this paper, we seek to address a key drawback of post-hoc concept-removal methods. Our contribution is to introduce SPLINCE, a projection that (similar to LEACE or SAL) prevents any linear classifier from predicting a concept, while also preserving the covariance with a task of interest. Mathematically, we construct an oblique projection that places the covariance between the representations and a protected attribute in its kernel, while maintaining the covariance between the representations and the main-task label in its range.
We prove that if a linear classifier is re-fitted after projection without regularization, any two projections that share the same kernel (i.e., that linearly erase the same subspace) will induce identical loss. In that sense, SPLINCE and previous methods such as LEACE (Belrose et al., 2023) (as well as more naive versions that do not explicitly aim to maintain the minimality of the projection) are all equivalent. For causal model interventions aiming to interpret its behavior—a situation where the underlying model is necessarily frozen—we argue that SPLINCE may perform a more surgical intervention, akin to minimizing the side effects of erasure on related concepts (e.g., removing gender bias while preserving grammatical gender). Empirically, we show that in a realistic classification setting, SPLINCE improves fairness in a highly challenging, imbalanced scenario, and removes stereotypes while maintaining correlated factual information.
@@ 2 Related work
Existing linear concept-removal methods: iterative nullspace projection (INLP, Ravfogel et al. (2020)) trains a linear classifier to predict a concept, and projects embeddings to the nullspace of the parameters of the linear classifier. This is repeated until the concept can no longer be predicted by the linear classifier. Relaxed Linear Adversarial Concept Erasure (RLACE, Ravfogel et al. (2022)) trains an orthogonal projection matrix such that a concept cannot be predicted by a linear classifier. These works were followed up by Least-squares Concept Erasure (LEACE, Belrose et al. (2023)), which ensures that no linear classifier can predict the concept (hereafter referred to as linear guardedness) while minimally altering the embeddings. Spectral attribute removal (SAL, Shao et al. (2023)) projects the embeddings orthogonal to the first $k$ eigenvectors of the covariance matrix $\mathrm{Cov}(\bm{x},\bm{z})$. Mean Projection (MP, (Haghighatkhah et al., 2022)) projects embeddings to the nullspace of the mean difference between embeddings with and without concepts. It is equivalent to SAL when the concept is binary. SAL and MP also guarantee linear guardedness (similar to LEACE), whereas other methods may or may not satisfy this criterion.
Linear concept-removal while preserving task-relevant information: previous work suggests that linear concept-removal methods remove task-relevant information in addition to the concept they seek to remove (Belinkov, 2022; Kumar et al., 2022; Guerner et al., 2025). In response, several alternatives have been proposed to address this issue, all removing a different linear subspace (Dev et al., 2021; Holstege et al., 2024; Bareeva et al., 2024; Shi et al., 2024). However, each of these approaches sacrifices linear guardedness in order to retain more task relevant information. An alternative approach is to explicitly optimize for fairness while maintaining task performance (Shen et al., 2021). However, this approach is more resource-intensive, as it cannot be applied to the frozen representations of a pretrained model. Moreover, because it modifies the original representations, it is unsuitable for scenarios where the intervention is intended to simulate causal experiments on the behavior of a pretrained LM, as discussed in section˜4.2.
@@ 3 Theory
3.1 SPLINCE: Ensuring linear guardedness while preserving task-relevant covariance
The LEACE projection ensures linear guardedness while minimizing the distortion of the features, but is oblivious of the main task of the model. A small expected norm squared may not optimally preserve information that is actually useful for the task at hand. Suppose now there is a random vector $\bm{y}\in\mathcal{Y}$ that represents the task-relevant information. Similar to the concept vector $\bm{z},$ it can be binary, one-hot encoded or continuous. We conjecture that task-relevant information in the features $\bm{x}$ is located in the directions that linearly covariate with $\bm{y}.$ In other words, it is located in the column space of the covariance matrix $\bm{\Sigma}_{\bm{x},\bm{y}}=\mathrm{Cov}(\bm{x},\bm{y})$. Indeed, removing this particular subspace completely prevents linear classification (Belrose et al., 2023).
In order to preserve this task-relevant information, we require the affine transformation $r(\bm{x})=\mathbf{P}\bm{x}+\bm{b}$ not only to produce features that are linearly guarded for $\bm{z},$ but also to leave the covariance between $\bm{x}$ and $\bm{y}$ invariant. For this approach we cast the name SPLINCE (Simultaneous Projection for LINear concept removal and Covariance prEservation), which can be seen as an extension of LEACE. It eliminates sensitive concepts from representations while exactly preserving their covariance with a target label. The SPLINCE optimization problem is formulated in Theorem 1 and its solution is given by equation 4, which is the main theoretical contribution of this paper.
@@ 3.1 SPLINCE: Ensuring linear guardedness while preserving task-relevant covariance
$\mathbf{P}^{\star}_{\mathrm{SPLINCE}}=\mathbf{W}^{+}\mathbf{V}(\mathbf{U}^{\mathrm{T}}\mathbf{V})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{W},$ (4)
@@
The proof of Theorem 1 is given in Appendix A.1. Compared to the LEACE optimization problem, the only difference is the additional condition of preserving task-relevant information, $\mathbf{P}\bm{\Sigma}_{\bm{x},\bm{y}}=\bm{\Sigma}_{\bm{x},\bm{y}},$ which we refer to as the range constraint. Similar to LEACE, $\mathbf{P}^{\star}_{\mathrm{SPLINCE}}$ is an oblique transformation with kernel $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}}).$ The difference lies in the range, which now contains $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ in order to fulfill the range constraint. Intuitively, the freedom that the LEACE optimization problem gives to the choice of the range is partially used to preserve the task-relevant information, i.e., the covariance between $\bm{x}$ and $\bm{y}.$ The remainder of the freedom is used to minimize the distortion caused by the affine transformation, leading to whitening and unwhitening, similar to LEACE.
The main assumption of Theorem 1 is that $\mathcal{U}^{\perp}\cap\mathrm{colsp}(\mathbf{W}\bm{\Sigma}_{\bm{x},\bm{y}})=\{\mathbf{0}\}$, which is satisfied as long as the subspaces spanned by $\mathrm{Cov}(\bm{x},\bm{z})$ and $\mathrm{Cov}(\bm{x},\bm{y})$ do not perfectly overlap. For the case where $\bm{z}$ and $\bm{y}$ are binary variables, this assumption is equivalent to requiring that the covariance vectors $\mathrm{Cov}(\bm{x},\bm{z})$ and $\mathrm{Cov}(\bm{x},\bm{y})$ are linearly independent (i.e., not proportional).
We note that, perhaps counter-intuitively, SPLINCE is not necessarily equivalent to LEACE if $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}})$ and $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ are orthogonal subspaces. Only if those subspaces are orthogonal after whitening, SPLINCE and LEACE are equivalent. In that case the orthogonal projection of LEACE then already contains the task-relevant directions $\mathrm{colsp}(\mathbf{W}\bm{\Sigma}_{\bm{x},\bm{y}})$ in its range. We also note that, similar to LEACE (Belrose et al., 2023), in the case of non-centered data, i.e., $\mathbb{E}[\bm{x}]\neq\mathbf{0},$ the optimal affine transformation requires the addition of a constant $\bm{b}^{\star}_{\mathrm{SPLINCE}}=\mathbb{E}[\bm{x}]-\mathbf{P}^{\star}_{\mathrm{SPLINCE}}\mathbb{E}[\bm{x}].$ Finally, in Figure 1 we give a visual illustration of the steps of the projection matrix suggested by Theorem 1.
@@ 3.2 Last-layer linear concept removal with re-training
The first use case of SPLINCE we consider is linear concept removal applied to the embeddings of a DNN, after which a linear classifier is fitted on the transformed embeddings. This can be useful if it is demanded that a predictive model does not make use of sensitive concepts, like gender or race.
If we compare SPLINCE with other concept removal methods that guarantee linear guardedness, namely LEACE and SAL (Shao et al., 2023), we observe that they are all projections with the same kernel $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}}).$ They differ in the choice of the range. Interestingly, we find that the predictions of a linear classifier that is trained without regularization on the transformed embeddings are not affected by this choice of the range. In other words, all concept removal methods that ensure linear guardedness will lead to the same predictions after re-training a linear classifier without regularization. We verify this empirically in Appendix B.1.
This result is formalized in Theorem 3.2, in which we consider training a model $f(\bm{x};\bm{\theta})$ that only depends on the embeddings $\bm{x}$ and parameters $\bm{\theta}$ through their inner product. Examples of such models are linear and logistic regression, the latter typically being used as the linear classifier re-trained on the embeddings. Before fitting, a projection is applied to the embeddings. We consider two projections that have the same kernel, but different ranges. Theorem 3.2 shows that, in the case of a strictly convex loss function without regularization, both fitted models lead to the same predictions.
@@ 3.3 When does changing the range matter?
Theorem 3.2 provides a case where SPLINCE will lead to the same predictions as other concept removal methods that ensure linear guardedness (e.g., SAL and LEACE). Here, we identify two practical cases where applying projections with the same kernel and different ranges will typically lead to different predictions.
@@
    When re-training the last layer with regularization: if we include a regularization term (such as $\bm{\theta}||_{2}$ or $\bm{\theta}||_{1}$) in our loss function $\mathcal{L}$, then it no longer exclusively depends on the parameters via the inner product $\mathrm{\mathbf{X}}\bm{\theta}.$ This will generally lead to $\bm{x}_{\mathcal{A}}^{\mathrm{T}}\bm{\theta}_{\mathcal{A}}^{*}\neq\bm{x}_{\mathcal{B}}^{\mathrm{T}}\bm{\theta}_{\mathcal{B}}^{*}$ for any two projections with the same kernel and different ranges.
@@
    Since the projections considered in this paper are constructed without knowledge of the parameters, these will typically lie outside the orthogonal complements. Note that this use case is relevant for language modeling, where re-training the parameters of the last layer is typically not feasible in terms of computational resources and/or data availability.
For these cases we expect SPLINCE to outperform other methods that ensure linear guardedness, as it is designed to preserve task-relevant information. We empirically investigate this in the next section. [^1] See this link for our code for the experiments, as well as an implementation of SPLINCE.
@@ 4 Experiments
This section is structured as follows. We start by investigating classification tasks when the last layer of the model is re-fitted with regularization. Then, in Section 4.2, we investigate language modeling, when the last layer of the language model (LM) is not re-trained post-projection. Finally, in order to qualitatively assess the effect of the different projections, we apply SPLINCE to black and white image data in Section 4.3. Across experiments, we compare SPLINCE to two other projections: LEACE and SAL (see Section 2). We focus on these projections since, similar to SPLINCE, they guarantee linear guardedness with regards to a concept, and only differ in choice of range. Furthermore, LEACE and SAL have been shown to outperform other existing concept-removal methods such as INLP and RLACE, which may or may not satisfy linear guardedness.
@@ 4.1 Classification where the last layer is re-trained with regularization
Set-up of the experiment: we seek to investigate the impact of each projection as the correlation between the task of interest $(y_{\mathrm{prof}},y_{\mathrm{tox}})$ and the concept to remove $(z_{\mathrm{gender}},z_{\mathrm{lang}})$ becomes stronger. We expect that as the relationship becomes stronger, the difference between SPLINCE and other projections becomes greater. The reason is that stronger correlated labels have covariances with the embeddings that are typically more aligned. Removing the concept is then more likely to also remove information about the task of interest.
@@
Results: the results of the experiments for both datasets are given in Figure 2. We focus on overall accuracy as well as worst-group accuracy. Worst-group accuracy is defined as the lowest accuracy for all combinations of the task and concept. A low worst-group accuracy reflects that a model relies on the correlation between the task and concept in the training data (Sagawa et al., 2020). For both datasets, as the relationship between task and concept becomes stronger, SPLINCE outperforms the other projections in both accuracy and worst-group accuracy. As the correlation between the task and concept becomes stronger, SAL and LEACE remove a significant part of $\bm{\Sigma}_{\bm{x},\bm{y}},$ contrary to SPLINCE. This is illustrated for the Bias in Bios dataset in Figure 7 in Appendix B.2